OpenStudy (anonymous):

Just checking:

OpenStudy (anonymous):

When solving the system x + 2y = 9 and x - y = 6, which expression could be substituted for x in the second equation? A. 9 - x B. y - 9/2 C. 9 - 2y D. 1/2 y + 9

OpenStudy (anonymous):

I think it's B.

OpenStudy (anonymous):

x + 2y = 9 -2y -2y ------------ x = 9 - 2y

OpenStudy (anonymous):

So it is B?

OpenStudy (anonymous):

No, its C

OpenStudy (anonymous):

9 - 2y

OpenStudy (anonymous):

oh okay I misread it ;) thnk u Tyler